# Fourier Dimension in Duffin–Schaeffer Conjecture

### 15 Apr 2026 v1

@@ Abstract
Let $\psi\colon\mathbb{N}\to[0,\frac{1}{2})$ be a positive function, and let $\theta\colon\mathbb{N}\to\mathbb{R}$ be a real-valued function. Let $\{A_{q}\}$ and $\{B_{q}\}$ be sequences of integers such that $\gcd(A_{q},B_{q})=1$ and $B_{q}>0$ for all $q$. Define $W^{\ast}(\psi,\theta)$ to be the set of $x\in[0,1]$ for which
$\left|x-\frac{p+\theta(q)}{q}\right|<\frac{\psi(q)}{q}\text{ for infinitely many }(p,q)\in\mathbb{Z}\times\mathbb{N}\text{ with }\gcd(B_{q}p+A_{q},q)=1.$
In this paper, we determine the Fourier dimension of $W^{*}(\psi,\theta)$. Our result not only recovers the classical theorems of Kaufman and Bluhm (concerning the homogeneous case $\psi(q)=q^{-\tau}$, $\tau\geq 1$) but also furnishes a complete inhomogeneous generalization. In addition, it affirmatively resolves the coprime formulation of the Chen and Xiong conjecture.
@@
The non‑monotonic case is more involved. In the homogeneous setting $(\bm{\theta}=\bm{0})$, Beresnevich and Velani [4] removed the monotonicity condition when $nm\geq 2.$ However, monotonicity is necessary in the case $(m,n)=(1,1),$ as demonstrated by the counterexample of Duffin and Schaeffer [16]. Further, consider the set
@@
For details and a comprehensive overview of the Hausdorff dimension of $W(n,m;\psi,{\bm{\theta}})$ in higher dimensions, see the papers [45, 5] and the references therein.
@@
where $\langle{\bm{t}},\mathbf{x}\rangle=t_{1}x_{1}+\cdots+t_{d}x_{d}.$ Loosely speaking, Fourier decay reflects the “smoothness" or “randomness" of a measure: it indicates that the measure is spread out in such a way that coherent long‑range correlations are dampened. In harmonic analysis and geometric measure theory, this decay is intimately connected to the fractal and dimensional properties of the measure’s support. Indeed, a classical result of Frostman states that for $E\subset\mathbb{R}^{d}$,
@@
  • Let $Q\subseteq\mathbb{N}$ be infinite, and define the associated approximable set
@@
    Specially, $W_{\mathbb{N}}(1,1;\psi,\theta)=W(1,1;\psi,\theta).$ Cai and Hambrook [8, Theorem 1.4.1] proved that under the convergence condition $\sum_{q\in Q}\psi(q)<\infty$,
@@
Let $\psi\colon\mathbb{N}\to[0,\tfrac{1}{2})$ be a positive function and let $\theta\colon\mathbb{N}\to\mathbb{R}$ be an arbitrary function. Let $\{A_{q}\}$ and $\{B_{q}\}$ be sequences of integers such that $\gcd(A_{q},B_{q})=1$ and $B_{q}>0$ for all $q$. Define $W^{\ast}(\psi,\theta)$ to be the set of $x\in[0,1]$ such that
$\left|x-\frac{p+\theta(q)}{q}\right|<\frac{\psi(q)}{q}\text{ for i.m. }(p,q)\in\mathbb{Z}\times\mathbb{N}\text{ with }\gcd(B_{q}p+A_{q},q)=1.$
Strictly speaking, the set ${W}^{\ast}(\psi,\theta)$ depends on the sequences $\{A_{q},B_{q}\}$. Since these sequences are fixed throughout our study, we omit this dependence from the notation. Thanks to the limsup set structure of $W^{*}(\psi,\theta)$, it follows from [2, Proposition 1] and the first Borel–Cantelli lemma that the Lebesgue measure of $W^{*}(\psi,\theta)$ vanishes provided that
$\sum_{q\in\mathbb{N}}\psi(q)\prod_{{p\mid q,\penalty 10000\ p\nmid B_{q}}}\bigl(1-p^{-1}\bigr)<\infty,$ (3)
@@
Let $\psi\colon\mathbb{N}\to[0,\frac{1}{2})$ be an arbitrary function satisfying the condition (3). Then the Fourier dimension of $W^{*}(\psi,\theta)$ is
$\dim_{{\rm F}}W^{\ast}(\psi,\theta)=\min\{2s(\psi),1\}.$
@@
    Then $\left|x-({p+\theta})/{q}\right|$ can be well approximated by $\left|x-({B_{q}p+A_{q}})/({B_{q}q})\right|$. Therefore, analogous to the homogeneous Duffin-Schaeffer conjecture, the restriction
@@ 2 Proof of Theorem 1
To begin, we recall the limsup set structure of $W^{\ast}(\psi,\theta)$. For $q\in\mathbb{N},$ define
@@
[…]

### 16 Apr 2026 v2

# Fourier Dimension in Inhomogeneous Duffin–Schaeffer Conjecture

@@ Abstract
Let $Q\subseteq\mathbb{N}$ be a subset, and let $\psi\colon\mathbb{N}\to[0,\tfrac{1}{2})$, $\theta\colon\mathbb{N}\to\mathbb{R}$ be functions. Let $\{A_{q}\}$ and $\{B_{q}\}$ be sequences of integers such that $\gcd(A_{q},B_{q})=1$ and $B_{q}>0$ for all $q$. Define $W_{Q}^{\ast}(\psi,\theta)$ to be the set of $x\in[0,1]$ for which
$\left|x-\frac{p+\theta(q)}{q}\right|<\frac{\psi(q)}{q}$
holds for infinitely many $(p,q)\in\mathbb{Z}\times Q$ with $\gcd(B_{q}p+A_{q},q)=1$.
In this paper, we determine the Fourier dimension of $W_{Q}^{\ast}(\psi,\theta)$. Our result not only recovers the classical theorems of Kaufman and Bluhm (concerning the homogeneous case $\psi(q)=q^{-\tau}$ with $\tau\geq 1$) and the one-dimensional version of a result by Cai and Hambrook on the inhomogeneous approximable set, but also provides a complete inhomogeneous generalization. Moreover, it gives an affirmative answer to the coprime formulation of the Chen–Xiong conjecture.
@@
The non-monotonic case is more involved. In the homogeneous setting $(\bm{\theta}=\bm{0})$, Beresnevich and Velani [4] removed the monotonicity condition when $nm\geq 2.$ However, monotonicity is necessary in the case $(m,n)=(1,1),$ as demonstrated by the counterexample of Duffin and Schaeffer [16]. Further, consider the set
@@
For details and a comprehensive overview of the Hausdorff dimension of $W(n,m;\psi,{\bm{\theta}})$ in higher dimensions, see [45, 5] and the references therein.
@@
where $\langle{\bm{t}},\mathbf{x}\rangle=t_{1}x_{1}+\cdots+t_{d}x_{d}.$ Loosely speaking, Fourier decay reflects the “smoothness" or “randomness" of a measure: it indicates that the measure is spread out in such a way that coherent long-range correlations are dampened. In harmonic analysis and geometric measure theory, this decay is intimately connected to the fractal and dimensional properties of the measure’s support. Indeed, a classical result of Frostman states that for $E\subset\mathbb{R}^{d}$,
@@
  • Let $Q\subseteq\mathbb{N}$ be a subset, and define the associated approximable set
@@
    Specially, $W_{\mathbb{N}}(1,1;\psi,\theta)=W(1,1;\psi,\theta).$ Cai and Hambrook [8, Theorem 1.4.1] [^3] Cai and Hambrook studied the mutli-dimensional generalization of $W_{Q}(1,1;\psi,\theta)$ proved that under the convergence condition $\sum_{q\in Q}\psi(q)<\infty$,
@@
Let $\{A_{q}\}$ and $\{B_{q}\}$ be sequences of integers such that $\gcd(A_{q},B_{q})=1$ and $B_{q}>0$ for all $q$. Let $Q\subseteq\mathbb{N}$ be a subset, let $\psi\colon Q\to[0,\tfrac{1}{2})$ be a positive function, and let $\theta\colon Q\to\mathbb{R}$ be an arbitrary function. Define $W_{Q}^{\ast}(\psi,\theta)$ to be the set of $x\in[0,1]$ such that
$\left|x-\frac{p+\theta(q)}{q}\right|<\frac{\psi(q)}{q}\text{ for i.m. }(p,q)\in\mathbb{Z}\times Q\text{ with }\gcd(B_{q}p+A_{q},q)=1.$
Strictly speaking, the set ${W}_{Q}^{\ast}(\psi,\theta)$ depends on the sequences $\{A_{q},B_{q}\}$. Since these sequences are fixed throughout our study, we omit this dependence from the notation. Thanks to the limsup set structure of $W_{Q}^{*}(\psi,\theta)$, it follows from [2, Proposition 1] and the first Borel–Cantelli lemma that the Lebesgue measure of $W^{*}(\psi,\theta)$ vanishes provided that
$\sum_{q\in Q}\psi(q)\prod_{{p\mid q,\penalty 10000\ p\nmid B_{q}}}\bigl(1-p^{-1}\bigr)<\infty,$ (3)
@@
Let $Q\subset\mathbb{N}$, and let $\psi\colon Q\to[0,\frac{1}{2})$ be a function satisfying the condition (3). Then the Fourier dimension of $W_{Q}^{*}(\psi,\theta)$ is
$\dim_{{\rm F}}W_{Q}^{\ast}(\psi,\theta)=\min\{2s(Q,\psi),1\}.$
When $Q=\mathbb{N}$, we abbreviate $W_{\mathbb{N}}^{\ast}(\psi,\theta)$ as $W^{\ast}(\psi,\theta)$. A direct corollary of Theorem 1 obtained by taking $Q=\mathbb{N}$ reads as follows.
Corollary 2.
Let $\psi\colon\mathbb{N}\to[0,1/2)$ be a function satisfying condition (3) with $Q$ replaced by $\mathbb{N}$. Then
$\dim_{\mathrm{F}}W^{\ast}(\psi,\theta)=\min\{2s(\psi),1\}.$
Remark 2. Although Corollary 2 is a special case of Theorem 1, they are essentially equivalent. Indeed, assuming Corollary 2 holds, we prove Theorem 1 as follows.
We extend the function $\psi:Q\to[0,1/2)$ to $\overline{\psi}:\mathbb{N}\to[0,1/2)$ by setting
$\overline{\psi}(q)=\begin{cases}\psi(q)&\text{if }q\in Q,\\[4.0pt] 0&\text{if }q\notin Q.\end{cases}$
(The value of $\theta$ on $\mathbb{N}\setminus Q$ may be defined arbitrarily.) One then readily verifies that
$W_{Q}^{\ast}(\psi,\theta)=W^{\ast}(\overline{\psi},\theta)\qquad\text{and}\qquad s(Q,\psi)=s(\overline{\psi}).$
@@
    Then $\left|x-({p+\theta})/{q}\right|$ can be well approximated by $\left|x-({B_{q}p+A_{q}})/({B_{q}q})\right|$. Therefore, analogous to the homogeneous Duffin–Schaeffer conjecture, the restriction
@@
  • In the special case where $A_{q}=1$ and $B_{q}=q$ for all $q\in\mathbb{N}$, our result also recovers Theorem 1.4.1 of [8] in the one-dimensional setting.
@@ 2 Proof of Theorem 1
As claimed before, it suffices to prove Corollary 2. To begin, we recall the limsup set structure of $W^{\ast}(\psi,\theta)$. For $q\in\mathbb{N},$ define
@@
$\widehat{\mathbf{I}}_{A_{p,q}}(0)=\frac{2\psi(q)}{q},\quad|\widehat{\mathbf{I}}_{A_{p,q}}(k)|\ll\frac{1}{|k|}\penalty 10000\ \penalty 10000\ \big(k\in\mathbb{Z}\setminus\{0\}\big).$
@@ Acknowledgements
Declarations
Conflict of interest On behalf of all authors, the corresponding author states that there is no conflict of interest.